# Variational Search Distributions

### v2

@@ D.5 Performance analysis
Appendix E VSD with neural network CPEs
@@
The result in E.1 suggests that minimizing the cross-entropy loss will lead us to minimize the MSE loss as well, since the latter is upper bounded by the former. This result provides us with theoretical justification to derive convergence results based on the MSE loss, which has been better analyzed in the NN literature (Jacot et al., 2018; Lee et al., 2019), as a proxy to establish convergence guarantees for the CE-based VSD setting.
E.1 Linear approximations via the neural tangent kernel
Let $\pi^{*}$ denote the unknown true classifier, i.e., $\pi(\mathbf{x}):=p({y>\tau}|{\mathbf{x},{f\!\centerdot}})$, for $\mathbf{x}\in\mathcal{X}$. In the following, we will assume that $\pi^{*}$ is an unknown, fixed element of a reproducing kernel Hilbert space (RKHS) associated with a given kernel (Schölkopf & Smola, 2001). In the case of infinite-width neural networks, we know that under certain assumptions the NN trained via gradient descent under the MSE loss will asymptotically converge to a kernel ridge regression solution whose kernel is given by the neural tangent kernel (NTK, Jacot et al. (2018)). This asymptotic solution is equivalent to the posterior mean of a Gaussian process that assumes no observation noise. For a finite amount of training steps with a non-infinitesimal learning rate, however, the literature has shown that gradient-based training provides a form of implicit regularization. In that case, we recover a regularized kernel ridge regression solution which can be robust to label noise (Hu et al., 2020) for us to use. Lastly, although our analysis will be based on NTK results, the approximation error between the infinite-width and the finite-width NN vanishes with the square root of the network width for most popular NN architectures (Liu et al., 2020). Therefore, we can assume that these approximation guarantees will remain useful for wide-enough, finite-width NN models.
@@ E.1 Linear approximations via the neural tangent kernel
Several results in the literature have shown that training overparameterized neural networks via gradient descent provides a form of implicit regularization on the learned model (Fleming, 1990; Yao et al., 2007; Soudry et al., 2018; Barrett & Dherin, 2021), with some of the same behavior extending to the stochastic gradient setting (Smith et al., 2021). In particular, Fleming (1990) showed a direct equivalence between an early stopped gradient-descent linear model and the solution of a regularized least-squares problem with a penalty on the parameters vector Euclidean norm. Therefore, as wide deep neural networks behave as linear models in the infinite-width limit when trained via gradient descent (Jacot et al., 2018; Lee et al., 2019), it is reasonable to model our least-squares problem with the MSE loss via its regularized version:
$\hat{\pi}_{N}\in\operatorname*{arg\,\!min}_{\pi\in{\mathcal{{F}}}_{k}}\sum_{i=1}^{N}(\pi(\mathbf{x}_{i})-z_{i})^{2}+\rho\lVert\pi\rVert_{k}^{2},$ (90)
where $\rho>0$ is a possibly unknown regularization factor, ${\mathcal{{F}}}_{k}$ denotes the RKHS associated with the neural tangent kernel $k$ for the given CPE NN architecture, and $\lVert\cdot\rVert_{k}$ denotes the RKHS norm. The problem above is equivalent to regularized kernel ridge regression (Shawe-Taylor & Cristianini, 2004), whose solution is given by:
$\hat{\pi}_{N}(\mathbf{x})=\pi_{0}(\mathbf{x})+{\boldsymbol{\mathbf{k}}}_{N}(\mathbf{x})^{\top}({\boldsymbol{\mathbf{{K}}}}_{N}+\rho\mathbf{I})^{-1}({\boldsymbol{\mathbf{z}}}_{N}-\pi_{0}(\mathcal{X}_{N})),$ (91)
@@
where ${\boldsymbol{\mathbf{k}}}_{N}$ and ${\boldsymbol{\mathbf{{K}}}}_{N}$ are defined in the same way as for the GP case, $\pi_{0}$ denotes the untrained NN classifier at initialization, and $\pi_{0}(\mathcal{X}_{N}):=[\pi_{0}(\mathbf{x}_{i})]_{i=1}^{N}$. For our analysis, we may assume that $\pi_{0}=0$ at times, noting that the least-squares problem can always be solved for the residuals $z-\pi_{0}(\mathbf{x})$ and then have $\pi_{0}$ added back to the solution. We refer the reader to Jacot et al. (2018) for further discussion on the effect of the network initialization.
@@
For fully connected, convolutional or residual networks equipped with smooth activation functions (e.g., sigmoid or $\tanh$), Liu et al. (2020) showed that the approximation error between the linear model and the finite-width NN is $\widetilde{{\mathcal{{O}}}}(m^{-1/2})$, where $m$ denotes the minimum layer width, and the $\widetilde{{\mathcal{{O}}}}$ notation corresponds to the ${\mathcal{{O}}}$-notation with logarithmic factors suppressed. NTK results for other activation functions, e.g., ReLU (Chen & Xu, 2021), and different neural network architectures, such as multi-head attention (Hron et al., 2020), are also available in the literature.
E.2 Assumptions
In the following, we present a series of mild technical assumptions needed for our theoretical analysis of NN-based CPEs. Firstly, we assume a bounded NTK.
Assumption E.1.
The NTK $k$ corresponding to the network architecture in $\pi_{\theta}$ is bounded:
$\sup_{\mathbf{x}\in\mathcal{X}}k(\mathbf{x},\mathbf{x})<\infty\,,$ (92)
for some constant $b_{k}>0$.
We will also assume that the threshold is fixed to simplify the analysis. However, our results should still be applicable to the time-varying threshold setting after minor adjustments.
Assumption E.2.
The threshold is fixed, i.e., $\tau_{t}=\tau\in$, for all $t\geq 1$.
The following assumption on label noise should always hold for Bernoulli random variables (Boucheron et al., 2013). Any upper bound on the sub-Gaussian parameter should suffice for the analysis (e.g., $\sigma_{z}\leq 1$ for Bernoulli variables).
Assumption E.3.
For all $t\in\mathbb{N}$ and all $\mathbf{x}\in\mathcal{X}$, label noise $\zeta=\mathbbm{1}[{y>\tau}]-\pi^{*}(\mathbf{x})$, with $y\sim p({y}|{\mathbf{x},{f\!\centerdot}})$, is $\sigma_{z}$-sub-Gaussian:
$\forall a\in,\quad\mathbb{E}\left[\exp\left(a\zeta\right)\right]\leq\exp\left(\frac{a^{2}\sigma_{z}^{2}}{2}\right),$ (93)
for some $\sigma_{z}\geq 0$.
For this analysis, we mainly assume that the true classifier $\pi^{*}(\mathbf{x})=p({y>\tau}|{\mathbf{x},{f\!\centerdot}})$ is a fixed, though unknown, element of the RKHS given by the NTK, which is formalized by the following assumption.
Assumption E.4.
@@ E.2 Assumptions
For a rich enough RKHS, such assumption is mild, especially given that most popular NN architectures possess universal approximation guarantees (Hornik et al., 1989). Finally, the next assumption ensures enough sampling asymptotically over the domain $\mathcal{X}$, which we still assume is finite.
Assumption E.5.
@@
$\exists b>0:\quad\forall t\in\mathbb{N},\quad q({\mathbf{x}}|{\phi_{t}})\geq b\,.$ (95)
@@
E.3 Approximation error for NN-based CPEs

### v4

@@ E.5 Performance analysis
Appendix F VSD with neural network CPEs
@@
The result in F.1 suggests that minimizing the cross-entropy loss will lead us to minimize the MSE loss as well, since the latter is upper bounded by the former. This result provides us with theoretical justification to derive convergence results based on the MSE loss, which has been better analyzed in the NN literature (Jacot et al., 2018; Lee et al., 2019), as a proxy to establish convergence guarantees for the CE-based VSD setting.
F.1 Linear approximations via the neural tangent kernel
For this analysis, we will follow a frequentist setting. Namely, let $\pi^{*}$ denote the unknown true classifier, i.e., $\pi(\mathbf{x}):=p({y>\tau}|{\mathbf{x},{f\!\centerdot}})$, for $\mathbf{x}\in\mathcal{X}$. We assume that $\pi^{*}$ is an unknown, fixed element of a reproducing kernel Hilbert space (RKHS) associated with a given kernel (Schölkopf & Smola, 2001). In the case of infinite-width neural networks, we know that under certain assumptions the NN trained via gradient descent under the MSE loss will asymptotically converge to a kernel ridge regression solution whose kernel is given by the neural tangent kernel (NTK, Jacot et al., 2018). This asymptotic solution is equivalent to the posterior mean of a Gaussian process that assumes no observation noise. However, for a finite number of training steps $s<\infty$, the literature has shown that gradient-based training provides a form of implicit regularization, which we use to ensure robustness to label noise. Moreover, although our analysis will be based on the NTK, the approximation error between the infinite-width and the finite-width NN vanishes with the square root of the network width for most popular NN architectures (Liu et al., 2020). Therefore, we can assume that these approximation guarantees will remain useful for wide-enough, finite-width NN models.
Reproducing kernel Hilbert spaces.
The RKHS ${\mathcal{{F}}}_{k}$ associated with a positive-semidefinite kernel $k:\mathcal{X}\times\mathcal{X}\to$ is a Hilbert space of functions over $\mathcal{X}$ with an inner product $\langle\cdot,\cdot\rangle_{k}$ and corresponding norm $\lVert\cdot\rVert_{k}:=\sqrt{\langle\cdot,\cdot\rangle}_{k}$ such that, for every $\pi\in{\mathcal{{F}}}_{k}$, the reproducing property $\pi(\mathbf{x})=\langle\pi,k(\cdot,\mathbf{x})\rangle_{k}$ holds for all $\mathbf{x}\in\mathcal{X}$ (Schölkopf & Smola, 2001).
@@ F.1 Linear approximations via the neural tangent kernel
Several results in the literature have shown that training overparameterized neural networks via gradient descent provides a form of implicit regularization on the learned model (Fleming, 1990; Yao et al., 2007; Soudry et al., 2018; Barrett & Dherin, 2021), with some of the same behavior extending to the stochastic gradient setting (Smith et al., 2021). In earlier works, Fleming (1990) showed a direct equivalence between an early stopped gradient-descent linear model and the solution of a regularized least-squares problem with a penalty on the parameters vector Euclidean norm. In the NTK regime, the network output predictions at iteration $s\in\mathbb{N}$ of gradient descent are given by (Lee et al., 2019):
$\hat{\pi}_{N}(\mathbf{x})=\pi_{0}(\mathbf{x})+{\boldsymbol{\mathbf{k}}}_{N}(\mathbf{x})^{\top}{\boldsymbol{\mathbf{{K}}}}_{N}^{-1}(\mathbf{I}-e^{-\nu s{\boldsymbol{\mathbf{{K}}}}_{N}})({\boldsymbol{\mathbf{z}}}_{N}-\pi_{0}(\mathcal{X}_{N}))\,,\quad\mathbf{x}\in\mathcal{X}\,,$ (90)
where $\pi_{0}$ represents the network's initialization, $\nu>0$ denotes the learning rate, the kernel $k$ corresponds to the NTK associated with the given architecture, and the data is represented by $\mathcal{X}_{N}:=\{\mathbf{x}_{i}\}_{i=1}^{N}\subset\mathcal{X}$ and ${\boldsymbol{\mathbf{z}}}_{N}:=[z_{i}]_{i=1}^{N}\in\{0,1\}^{N}$. Rearranging terms and performing basic algebraic manipulations, the equation above can be shown to be equivalent to:
$\hat{\pi}_{N}(\mathbf{x})=\pi_{0}(\mathbf{x})+{\boldsymbol{\mathbf{k}}}_{N}(\mathbf{x})^{\top}({\boldsymbol{\mathbf{{K}}}}_{N}+{\boldsymbol{\mathbf{{\Sigma}}}}_{N})^{-1}({\boldsymbol{\mathbf{z}}}_{N}-\pi_{0}(\mathcal{X}_{N}))\,,\quad\mathbf{x}\in\mathcal{X}\,,$ (91)
where ${\boldsymbol{\mathbf{{\Sigma}}}}_{N}:={\boldsymbol{\mathbf{{K}}}}_{N}(e^{\nu s{\boldsymbol{\mathbf{{K}}}}_{N}}-\mathbf{I})^{-1}$ corresponds to a data-dependent regularization matrix. The above is equivalent to the solution of the a regularized least-squares problem, as we show below.
Lemma F.1.
Assume $k:\mathcal{X}\times\mathcal{X}\to$ is positive definite and $\pi_{0}=0$. Then Equation 91 solves the following regularized least-squares problem:
$\hat{\pi}_{N}\in\operatorname*{arg\,\!min}_{\pi\in{\mathcal{{F}}}_{k}}\sum_{i=1}^{N}(\pi(\mathbf{x}_{i})-z_{i})^{2}+\lVert{\boldsymbol{\mathbf{{R}}}}_{N}^{1/2}\pi\rVert_{k}^{2}\,,$ (92)
where ${\boldsymbol{\mathbf{{R}}}}_{N}:={\boldsymbol{\mathbf{{\Phi}}}}_{N}(e^{-\nu s{\boldsymbol{\mathbf{{K}}}}_{N}}-\mathbf{I})^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{N}^{\top}$, ${\boldsymbol{\mathbf{{\Phi}}}}_{N}:=[\varphi(\mathbf{x}_{1}),\dots,\varphi(\mathbf{x}_{N})]$, and $\varphi(\mathbf{x}):=k(\cdot,\mathbf{x})\in{\mathcal{{F}}}_{k}$ represents the kernel's canonical feature map, for $\mathbf{x}\in\mathcal{X}$.
Proof.
The least-squares loss can be rewritten as:
$\begin{split}\ell_{N}(\pi)&:=\sum_{i=1}^{N}(\pi(\mathbf{x}_{i})-z_{i})^{2}+\lVert\pi\rVert_{{\boldsymbol{\mathbf{{R}}}}_{N}}^{2}=\lVert{\boldsymbol{\mathbf{{\Phi}}}}_{N}^{\top}\pi-{\boldsymbol{\mathbf{z}}}_{N}\rVert_{2}^{2}+\lVert{\boldsymbol{\mathbf{{R}}}}_{N}^{1/2}\pi\rVert_{k}^{2}\,,\end{split}$ (93)
where $\lVert\cdot\rVert_{2}$ denotes the Euclidean norm of a vector. Taking the functional gradient with respect to $\pi\in{\mathcal{{F}}}_{k}$ and equating it to zero, we have that an optimal solution $\hat{\pi}$ satisfies:
$\begin{split}\nabla\ell_{N}(\hat{\pi})&=2{\boldsymbol{\mathbf{{\Phi}}}}_{N}({\boldsymbol{\mathbf{{\Phi}}}}_{N}^{\top}\hat{\pi}-{\boldsymbol{\mathbf{z}}}_{N})+2{\boldsymbol{\mathbf{{R}}}}_{N}\hat{\pi}=0\,.\end{split}$ (94)
The minimum-norm solution is then given by:
$\hat{\pi}=({\boldsymbol{\mathbf{{\Phi}}}}_{N}{\boldsymbol{\mathbf{{\Phi}}}}_{N}^{\top}+{\boldsymbol{\mathbf{{R}}}}_{N})^{+}{\boldsymbol{\mathbf{{\Phi}}}}_{N}{\boldsymbol{\mathbf{z}}}_{N}\,,$ (95)
where ${\boldsymbol{\mathbf{{A}}}}^{+}$ denotes the Moore-Penrose pseudo-inverse of an operator ${\boldsymbol{\mathbf{{A}}}}:{\mathcal{{F}}}_{k}\to{\mathcal{{F}}}_{k}$.
Let ${\boldsymbol{\mathbf{{\Phi}}}}_{N}={\boldsymbol{\mathbf{{U}}}}_{N}{\boldsymbol{\mathbf{{S}}}}_{N}{\boldsymbol{\mathbf{{V}}}}_{N}^{\top}$ represent the singular value decomposition of ${\boldsymbol{\mathbf{{\Phi}}}}_{N}$ in the RKHS (Mollenhauer et al., 2020), where ${\boldsymbol{\mathbf{{U}}}}_{N}:=[u_{1},\dots,u_{N}]$, ${\boldsymbol{\mathbf{{V}}}}_{N}:=[{\boldsymbol{\mathbf{v}}}_{1},\dots,{\boldsymbol{\mathbf{v}}}_{N}]$, with $\{u_{i}\}_{i=1}^{N}\subset{\mathcal{{F}}}_{k}$ and $\{{\boldsymbol{\mathbf{v}}}_{i}\}_{i=1}^{N}\subset^{N}$ denoting the left and right singular vectors, respectively, and ${\boldsymbol{\mathbf{{S}}}}_{N}\in^{N\times N}$ corresponds to the diagonal matrix of singular values of ${\boldsymbol{\mathbf{{\Phi}}}}_{N}$. There are $N$ non-zero singular values, since $k$ is assumed to be positive definite, and we have the correspondence ${\boldsymbol{\mathbf{{K}}}}_{N}={\boldsymbol{\mathbf{{\Phi}}}}_{N}^{\top}{\boldsymbol{\mathbf{{\Phi}}}}_{N}={\boldsymbol{\mathbf{{V}}}}_{N}{\boldsymbol{\mathbf{{\Lambda}}}}_{N}{\boldsymbol{\mathbf{{V}}}}_{N}^{\top}$ with ${\boldsymbol{\mathbf{{\Lambda}}}}_{N}={\boldsymbol{\mathbf{{S}}}}_{N}^{2}$ representing the diagonal matrix of eigenvalues of ${\boldsymbol{\mathbf{{K}}}}_{N}$, which is full-rank for a positive-definite kernel with distinct entries $\mathcal{X}_{N}:=\{\mathbf{x}_{i}\}_{i=1}^{N}\subset\mathcal{X}$. Applying the SVD to derive the pseudo-inverse in Equation 95 then yields:
$\begin{split}\hat{\pi}&=({\boldsymbol{\mathbf{{\Phi}}}}_{N}{\boldsymbol{\mathbf{{\Phi}}}}_{N}^{\top}+{\boldsymbol{\mathbf{{\Phi}}}}_{N}(e^{-\nu s{\boldsymbol{\mathbf{{K}}}}_{N}}-\mathbf{I})^{-1}{\boldsymbol{\mathbf{{\Phi}}}}_{N}^{\top})^{+}{\boldsymbol{\mathbf{{\Phi}}}}_{N}{\boldsymbol{\mathbf{z}}}_{N}\\ &=({\boldsymbol{\mathbf{{U}}}}_{N}{\boldsymbol{\mathbf{{\Lambda}}}}_{N}{\boldsymbol{\mathbf{{U}}}}_{N}^{\top}+{\boldsymbol{\mathbf{{U}}}}_{N}{\boldsymbol{\mathbf{{S}}}}_{N}{\boldsymbol{\mathbf{{V}}}}_{N}^{\top}(e^{-\nu s{\boldsymbol{\mathbf{{V}}}}_{N}{\boldsymbol{\mathbf{{\Lambda}}}}_{N}{\boldsymbol{\mathbf{{V}}}}_{N}^{\top}}-\mathbf{I})^{-1}{\boldsymbol{\mathbf{{V}}}}_{N}{\boldsymbol{\mathbf{{S}}}}_{N}{\boldsymbol{\mathbf{{U}}}}_{N}^{\top})^{+}{\boldsymbol{\mathbf{{U}}}}_{N}{\boldsymbol{\mathbf{{S}}}}_{N}{\boldsymbol{\mathbf{{V}}}}_{N}^{\top}{\boldsymbol{\mathbf{z}}}_{N}\\ &={\boldsymbol{\mathbf{{U}}}}_{N}({\boldsymbol{\mathbf{{\Lambda}}}}_{N}+{\boldsymbol{\mathbf{{\Lambda}}}}_{N}(e^{-\nu s{\boldsymbol{\mathbf{{\Lambda}}}}_{N}}-\mathbf{I})^{-1})^{-1}{\boldsymbol{\mathbf{{S}}}}_{N}{\boldsymbol{\mathbf{{V}}}}_{N}^{\top}{\boldsymbol{\mathbf{z}}}_{N}\\ &={\boldsymbol{\mathbf{{U}}}}_{N}{\boldsymbol{\mathbf{{S}}}}_{N}{\boldsymbol{\mathbf{{V}}}}_{N}^{\top}{\boldsymbol{\mathbf{{V}}}}_{N}({\boldsymbol{\mathbf{{\Lambda}}}}_{N}+{\boldsymbol{\mathbf{{\Lambda}}}}_{N}(e^{-\nu s{\boldsymbol{\mathbf{{\Lambda}}}}_{N}}-\mathbf{I})^{-1})^{-1}{\boldsymbol{\mathbf{{V}}}}_{N}^{\top}{\boldsymbol{\mathbf{z}}}_{N}\\ &={\boldsymbol{\mathbf{{\Phi}}}}_{N}({\boldsymbol{\mathbf{{K}}}}_{N}+{\boldsymbol{\mathbf{{K}}}}_{N}(e^{-\nu s{\boldsymbol{\mathbf{{K}}}}_{N}}-\mathbf{I})^{-1})^{-1}{\boldsymbol{\mathbf{z}}}_{N}\\ &={\boldsymbol{\mathbf{{\Phi}}}}_{N}({\boldsymbol{\mathbf{{K}}}}_{N}+{\boldsymbol{\mathbf{{\Sigma}}}}_{N})^{-1}{\boldsymbol{\mathbf{z}}}_{N}\,,\end{split}$ (96)
which concludes the proof. ∎
@@
For our analysis, we will assume that the classifier network is zero initialized with $\pi_{0}=0$, noting that the least-squares problem can always be solved for the residuals $z-\pi_{0}(\mathbf{x})$ and then have $\pi_{0}$ added back to the solution. We refer the reader to Lee et al. (2019) for further discussion on the effect of the network initialization.
@@
For fully connected, convolutional or residual networks equipped with smooth activation functions (e.g., sigmoid or $\tanh$), Liu et al. (2020) showed that the approximation error between the linear model and the finite-width NN is $\widetilde{{\mathcal{{O}}}}(m^{-1/2})$, where $m$ denotes the minimum layer width, and the $\widetilde{{\mathcal{{O}}}}$ notation corresponds to the ${\mathcal{{O}}}$-notation with logarithmic factors suppressed. NTK results for other activation functions and different neural network architectures, such as multi-head attention (Hron et al., 2020), are also available in the literature.
F.2 Assumptions
In the following, we present a series of mild technical assumptions needed for our theoretical analysis of NN-based CPEs. For this analysis, we mainly assume that the true classifier $\pi^{*}(\mathbf{x})=p({y>\tau}|{\mathbf{x},{f\!\centerdot}})$ is a fixed, though unknown, element of the RKHS ${\mathcal{{F}}}_{k}$ given by a bounded NTK $k$, which is formalized by the following two assumptions. As in the GP case, we assume batches of size $B=1$ to simplify the analysis.
Assumption F.1.
@@ F.2 Assumptions
For a rich enough RKHS, such assumption is mild, especially given that most popular NN architectures offer universal approximation guarantees (Hornik et al., 1989).
Assumption F.2.
The NTK $k$ corresponding to the network architecture in $\pi_{\theta}$ is positive definite and bounded in $\mathcal{X}$.
We will also assume that the threshold is fixed to simplify the analysis. However, our results should asymptotically hold for time-varying thresholds as long as the limit $\lim_{t\to\infty}\tau_{t}=\tau$ exists.
Assumption F.3.
The threshold is fixed, i.e., $\tau_{t}=\tau\in$, for all $t\geq 1$.
The following assumption on label noise should always hold for Bernoulli random variables (Boucheron et al., 2013). Any upper bound on the sub-Gaussian parameter should suffice for the analysis (e.g., $\sigma_{\zeta}\leq 1$ for Bernoulli variables).
Assumption F.4.
For all $t\in\mathbb{N}$ and all $\mathbf{x}\in\mathcal{X}$, label noise $\zeta=\mathds{1}[{y>\tau}]-\pi^{*}(\mathbf{x})$, with $y\sim p({y}|{\mathbf{x},{f\!\centerdot}})$, is $\sigma_{\zeta}$-sub-Gaussian:
$\forall a\in,\quad\mathbb{E}\left[\exp\left(a\zeta\right)\right]\leq\exp\left(\frac{a^{2}\sigma_{\zeta}^{2}}{2}\right),$ (98)
for some $\sigma_{\zeta}\geq 0$.
The next assumption ensures a sufficient amount of sampling is asymptotically achieved over the domain $\mathcal{X}$, which we still assume is finite.
Assumption F.5.
@@
$\exists b>0:\quad\forall t\in\mathbb{N},\quad q({\mathbf{x}}|{\phi_{t}})\geq b,\quad\forall\mathbf{x}\in\mathcal{X}\,.$ (99)
@@
Assumption F.6.
The learning rate $\nu_{t}$ at each round $t$ is such that:
$0<\nu_{*}\leq\nu_{t}\leq\frac{1}{\lambda_{\max}({\boldsymbol{\mathbf{{K}}}}_{t})}\,,\quad\forall t\in\mathbb{N}\,,$ (100)
for some $\nu_{*}>0$, where $\lambda_{\max}(\cdot)$ denotes the maximum eigenvalue of a matrix, and ${\boldsymbol{\mathbf{{K}}}}_{t}$ denotes the NTK matrix evaluated at the training points available at iteration $t\geq 1$.
This last assumption ensures that a gradient descent algorithm is convergent (Fleming, 1990), though we use it to bound the spectrum of the implicit regularization matrix ${\boldsymbol{\mathbf{{\Sigma}}}}_{t}$ after a finite number of training steps $s<\infty$ (a.k.a. early stopping), which is needed for our results. We highlight that, under mild assumptions on the data distribution, $\lambda_{\max}({\boldsymbol{\mathbf{{K}}}}_{t})\in{\mathcal{{O}}}(1)$ w.r.t. the number of data points (Murray et al., 2023), so that the bound in F.6 will not vanish. Therefore, such assumption is easily satisfied by maintaining a sufficiently small learning rate.
F.3 Approximation error for NN-based CPEs